What values for x must be excluded in the following fraction?
x-3/(4x-5)(x+1)

Found 2 solutions by fractalier, josmiceli:
Answer by fractalier(2101)

Any value that makes the denominator go to zero...that is if
4x - 5 = 0 or x + 1 = 0, we must exclude that value for x.
Thus x cannot be 5/4 or -1. We say these are excluded from the domain of x.

Answer by josmiceli(6797)

All values of x that make the denominator zero must be excluded.
x-3/(4x-5)(x+1)
If either (4x - 5) or (x+1) is zero, the fraction becomes infinitely large.
4x - 5 = 0
4x = 5
x = 5/4
x + 1 = 0
x = -1
So, the values of x = 5/4 and x = -1 are excluded